CONFORMATION-SPECIFIC UV and IR SPECTROSCOPY OF CONFORMATIONALLY CONSTRAINED α/γ PEPTIDE FOLDAMERS

\underline{RYOJI KUSAKA}, Department of Chemistry, Purdue University, West Lafayette, IN 47907, and Department of Chemistry, Graduate School of Science, Hiroshima University, Higashi-Hiroshima, 739-8526, Japan; DI ZHANG, PATRICK WALSH, JOSEPH GORD, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907; BRIAN F. FISHER, and SAMUEL H. GELLMAN, Department of Chemistry, University of Wisconsin, Madison, WI 53706.

Synthetic foldamers composed of heterogeneous backbones offer constructs for building unique secondary structures. α/γpeptides juxtapose the α-amino acid sub-units typical of proteins with γ-amino sub-units. Gellman and co-workers have developed efficient syntheses of α/γ-peptides that incorporate a cyclohexyl ring constraint at the γ^{3}-γ^{4} bond to limit backbone torsional mobility, and found that they form helices held together by C=O(i)···H-N(i+3) H-bonds composing 12-membered rings both in solution and in crystalline form. ^{a} This talk will present a detailed study of the singleconformation double-resonance UV and IR spectroscopy of Ac-Ala-γ_{ACHC}-NH-benzyl (αγ) and Ac-γ_{ACHC}-Ala-NH-benzyl (γα) capped peptides, in which γ_{ACHC} residues are constrained by a cis cyclohexyl ring at γ^{3}-γ^{4} bond with an ethyl group at γ^{2} position. The two αγ and γα peptides have three amide groups that are the minimum length necessary to form a 12-membered H-bond. Conformational assignments were made using the NH stretch, C=O stretch (amide I), and NH bend (amide II) regions of the IR spectrum with the aid of DFT calculations. The double-resonance UV and IR spectroscopy uncovered the presence of 6 conformers for αγ and 4 conformers for γα. In the two peptides, three of ten structures incorporate bifurcated double rings made of 12-membered C=O(1)···H-N(3) ring, which is the first stage of the 12-helix, and 7- or 9-membered C=O(1)···H-N(2) ring via nearest-neighbor interaction. The other seven structures are constructed based on 5-, 7-, and 9-membered nearest-neighbor H-bonds. The similarities and differences between structures observed for the two two αγ and γα peptides will be discussed.

aL. Guo, Y. G. Chi, A. M. Almeida, I. A. Guzei, B. K. Parker, and S. H. Gellman, J Am Chem Soc, 2009, 131, 16018-16020.